A 40-mH inductor has a resistance of 25 S2. What are its
reactance and impedance at the frequencies of 0 Hz (dc), 60
Hz, and 500 Hz?
1
Expert's answer
2020-12-21T11:01:58-0500

The reactance is:


"X_L = 2\pi fL"

where "f" is the frequency, and "L = 40mH = 0.04H" is the inductance of the inductor. Thus, for different frequencies obtain:


"X_L =2\pi\cdot 0\cdot 0.04 = 0\Omega\\
X_L =2\pi\cdot 60\cdot 0.04 \approx 15.1\Omega\\
X_L =2\pi\cdot 500\cdot 0.04 \approx 125.7\Omega"

The impedance is:


"Z = \sqrt{R^2 + X_L^2}"

where "R = 25\Omega" is the resistance of the inductor. Thus, for different "X_L" from the previous step obtain:


"Z = \sqrt{25^2 + 0^2} = 25\Omega\\
Z = \sqrt{25^2 + 15.1^2} \approx 29.2\Omega\\
Z = \sqrt{25^2 + 125.7^2} \approx 128.1\Omega"